### v1

# HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models

Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. Instead of a single model, our key insight is to take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions. Experimental results demonstrate that our approach is able to produce realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff can generate realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the supplementary material for video results. Darker color indicates later frames in the sequence. Best viewed in color.
1 Introduction
Text-driven synthesis of 3D human-object interactions (HOIs) aims to generate motions for both the human and object that form coherent and semantically meaningful interactions. It enables virtual humans to naturally interact with objects, which has a wide range of applications in AR/VR, video games, and filmmaking, etc.
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and objects to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [Guo_2022_CVPR, kit2016] or textual descriptions [bhatnagar22behave].
On one hand, recent methods [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, wu2022saga, taheri2021goal, Pi_2023_ICCV] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e. the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect the real-world complexity. For instance, they focus on grasping small objects [ghosh2023imos], provide the object motion as conditioning [li2023object], predict deterministic interactions between the human and the same object without the diversity [xu2023interdiff, razali2023action], consider only a small set of interactions (e.g. sit/lift [kulkarni2023nifty], sit/lie down [hassan_samp_2021], sit [jiang2022chairs, zhang2022couch, Pi_2023_ICCV], grasp [wu2022saga, taheri2021goal]), or investigate a single type of object (e.g. chair [jiang2022chairs, zhang2022couch]).
In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig. 1. Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [tevet2023human] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules.
To our best knowledge, HOI-Diff is the first text-driven approach capable of generating realistic motions for various HOIs and different types of objects during the interactions. For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [bhatnagar22behave] with text descriptions, which bridges the gap of no data for text-driven 3D HOIs generation. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [rombach2022high, saharia2022photorealistic, ramesh2021zero] and achieved impressive results. Recent work [zhang2022motiondiffuse, tevet2023human, chen2023executing, karunratanakul2023gmd, rempeluo2023tracepace, ahn2023can, barquero2023belfusion, chen2023humanmac, dabral2023mofusion, shafir2023human, sun2023towards, tian2023transfusion, wei2023understanding, zhang2023generating, zhang2023remodiffuse, zhang2023tedi, xie2023omnicontrol] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [huang2023diffusion, zhao2023synthesizing, wang2022towards, wang2022humanise]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, Pi_2023_ICCV] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [li2023object] proposes a conditional diffusion framework that can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [ghosh2023imos] synthesizes the full-body human along with the 3D object motions from textual inputs. Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands. InterDiff [xu2023interdiff] predicts whole-body interactions with dynamic objects. However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information. Note that the interaction type is deterministic, given the object and past motion information. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt.
Multi-Human Motion Generation. Recent works [shafir2023human, liang2023intergen, shafir2023human] incorporate human-to-human interactions into the motion diffusion process. They usually also adopt cross-attention to exchange information from different human motions. Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
Affordance Estimation. The affordance estimation on 3D point cloud is studied in [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive].  [kim2014semantic] formulates this task as the point cloud segmentation task from the extracted geometric features.  [kim2015interactive] voxelizes the point cloud and creates affordance maps with interactive manipulation.  [kokic2017affordance, iriondo2021affordance] focuses on the object grasping affordances and some specific applications.  [mo2022o2o] learns the affordance for object-object interactions.  [ngyen2023open] addressed the task of open-vocabulary affordance learning. Overall affordance learning is a very challenging task even the shape of point cloud is given. Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
3 Method
3.1 Overview
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [Guo_2022_CVPR] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models. Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
Diffusion Model for 3D HOI Generation. Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$***We use superscripts $h$ and $o$ to denote human and object squence, respectively. Without a superscript, it means the 3D HOI sequence, containing both ${\bm{x}}^{h}$ and ${\bm{x}}^{o}$. Subscript is used for the diffusion denoising step. learns the reverse diffusion process to generate clean data from a Gaussian noise ${{\bm{x}}_{t}}$ with $T$ consecutive denoising steps
$p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},\bm{c}):=\mathcal{N}({{\bm{x}}_{t-1}},\mathbf{\mu}_{\theta}({{\bm{x}}_{t}},t,\bm{c}),(1-\alpha_{t})\mathbf{I}),$ (1)
where $t$ is the denoising step. Following [tevet2023human], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$. We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [nichol2021improved]
${\bm{\mu}}_{t}=\frac{\sqrt{\alpha_{t-1}}\beta_{t}}{1-\alpha_{t}}{\bm{x}}_{0}+\frac{\sqrt{1-\beta_{t}}(1-\alpha_{t-1})}{1-\alpha_{t}}{{\bm{x}}_{t}},$ (2)
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [Ho2020-ew] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
Simply adopting the diffusion model described in Eq.(1) would impose huge burden to the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description. In this paper, we propose HOI-Diff for 3D HOIs generation, which disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules. Fig. 2 shows the overview of our proposed approach. We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent. As shown in Fig. 3, it consists of two Transformer models [vaswani2017attention], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [tevet2023human]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predicts clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [vaswani2017attention]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [tevet2023human]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in LABEL:ablation.
Figure 4: Illustration of APDM architecture for affordance estimation. Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean affordance estimation. The object point cloud and textual prompt are taken as conditional input.
3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the first problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [xu2023interdiff] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Fig. 4, the input to this model includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
Affordance estimation in 3D point clouds itself is a notably challenging problem [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive], especially in the context of 3D HOI generation involving textual prompt. To address the limited availability of training data, we propose a simplified approach that diverges from predicting point-wise contact labels. Our method is grounded in two key hypotheses. Firstly, we posit that interactions with an object involve a maximum of two contact points. Secondly, we identify eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the primary interacting parts in most HOI scenarios. These hypotheses streamline the complexity of the problem, making it more tractable without significantly compromising accuracy. For instance, this approach effectively models common interactions such as grasping an object with both hands, sitting actions involving the pelvis, or lifting with a single hand. Our empirical observations suggest that these simplifications are valid in a wide range of HOI applications, providing a balance between model complexity and real-world applicability.
Specifically, at each diffusion time step $n$ of APDMWe note that APDM and HOI-DM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy affordance data ${\bm{y}}_{n}$ includes human contact labels $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, two object contact positions $\bm{y}_{n}^{o}\in\mathbb{R}^{2\times 3}$. In addition, our APDM also generates a binary object state label ${\bm{y}}_{n}^{s}\in\{0,1\}$ indicating whether it remains static or moves during the interaction. The generated binary label can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. APDM works similar to the diffusion denoising process described in Eq.(1). Its output include clean affordance information ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [dhariwal2021diffusion] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Algorithm 1 Affordance-guided Interaction Correction
1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$
3:$K$ = 1
4:for all $t$ from $T$ to $1$ do
5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$
6:     Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$
7:     if $t=1$ then
8:         $K=100$
9:     end if
10:     for all $k$ from $K$ to $1$ do
11:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$
12:         ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$
13:     end for
14:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$
15:     $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$
16:end for
17:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can purturb the generated human and object motion at each diffusion step $t$ as in [xie2023omnicontrol, karunratanakul2023guided, kulkarni2023nifty],
$\mathbf{{\bm{\mu}}}^{h}_{t}$ $=\mathbf{{\bm{\mu}}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}^{h}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}),$ (3)
$\mathbf{{\bm{\mu}}}^{o}_{t}$ $=\mathbf{{\bm{\mu}}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{{\bm{\mu}}^{o}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}).$ (4)
Here $\tau_{1}$ and $\tau_{2}$ are different strengths to control the guidance for human and object motion, respectively. Due to the sparseness of object motion features, we assign a larger value to $\tau_{2}$ compared to $\tau_{1}$. This applies greater strength to perturb object motion, facilitating feasible corrections for contacting joints. During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step, as illustrated in Algorithm 1.
How can we define the objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$? We consider three terms here. First, in the generated 3D HOIs, the human and object should be close to each other on the contacting points. We therefore minimize the distance between human contact joints and object contact points
$G_{con}=\sum_{i\in\{1,2\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively. If there is just a single predicted contacting joint, $i=1$. $R(\cdot)$ converts the human joint’s local positions to global absolute locations, and $V(\cdot)$ obtains the object’s contact point sequence from the predicted mean of object pose ${\bm{\mu}}^{o}_{t}$.
Second, the generated motion of dynamic objects typically follows human movement. However, we observe that when the human interacts with a static object, such as sitting on a chair, the object appears slightly moved. To address this, we utilize the predicted object states ${\bm{y}}_{0}^{s}$ to immobilize the object’s movement in the generated samples if the state is static, ensuring that proper contact is established between the human and the static object. The objective is defined as
$\small{G_{sta}=\sum_{l=1}^{L}{\bm{y}}_{0}^{s}(l)\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
where ${\bm{\mu}}_{t}^{o}(l)$ denotes the object’s 6DoF pose in the $l$-th frame. $\bar{{\bm{\mu}}}_{t}^{o}=\frac{1}{L}\sum_{l}{\bm{\mu}}_{t}^{o}(l)$, which is the average of predicted means of the object’s pose.
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to space limit, we explain it in the supplementary material.
Finally, we combine all these goal functions to as the final objective
$G=G_{cont}+\alpha G_{sta}+\beta G_{smo},$ (7)
where $\alpha$ and $\beta$ are weights for balance.
4 Experiments
Table 1: Quantitative results on the BEHAVE [bhatnagar22behave] dataset. We compare our method with three baselines adapted from existing models and the real data. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
User Study $\uparrow$
Real 0.096 0.324 7.244 - - -
[…]
MDM* [tevet2023human] 3.169 0.206 7.606 0.448 0.184 1.20%
PriorMDM* [shafir2023human] 4.352 0.215 8.186 0.416 0.270 21.42%
Ours (Full) 1.632 0.222 7.596 0.363 0.190 77.38%
Table 2: Ablation studies of our model’s variants on the BEHAVE dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
Supplementary Material

### v2

@@
-HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models
+https://neu-vi.github.io/HOI-Diff/
 Abstract
 We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts.
-Instead of a single model, our key insight is to take a modular design and decompose the complex task into simpler sub-tasks.
-We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches.
+To this end, we take a modular design and decompose the complex task into simpler sub-tasks.
+We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches.
 We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt.
@@
 To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions.
-Experimental results demonstrate that our approach is able to produce realistic HOIs with various interactions and different types of objects.
-Figure 1: HOI-Diff can generate realistic motions for 3D human-object interactions given a text prompt and object geometry.
-Please see the supplementary material for video results.
+Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
+Figure 1: HOI-Diff generates realistic motions for 3D human-object interactions given a text prompt and object geometry.
+Please see the sup.
+mat.
+for video results.
 Darker color indicates later frames in the sequence.
@@
 The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges.
-First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and objects to avoid floating objects.
+First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects.
 Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair).
 Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored.
-Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [Guo_2022_CVPR, kit2016] or textual descriptions [bhatnagar22behave].
-On one hand, recent methods [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, wu2022saga, taheri2021goal, Pi_2023_ICCV] can synthesize realistic human motions for HOIs for static objects only.
+Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [11, 34] or textual descriptions [3].
+On one hand, recent methods [23, 17, 12, 42, 58, 52, 44, 33] can synthesize realistic human motions for HOIs for static objects only.
 They usually synthesize the motion in the last mile of interaction, i.e.
 the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them.
-On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect the real-world complexity.
-For instance, they focus on grasping small objects [ghosh2023imos], provide the object motion as conditioning [li2023object], predict deterministic interactions between the human and the same object without the diversity [xu2023interdiff, razali2023action], consider only a small set of interactions (e.g.
-sit/lift [kulkarni2023nifty], sit/lie down [hassan_samp_2021], sit [jiang2022chairs, zhang2022couch, Pi_2023_ICCV], grasp [wu2022saga, taheri2021goal]), or investigate a single type of object (e.g.
-chair [jiang2022chairs, zhang2022couch]).
+On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity.
+For instance, they focus on grasping small objects [10], provide the object motion as conditioning [25], predict deterministic interactions between the human and the same object without the diversity [54, 37], consider only a small set of interactions (e.g.
+sit/lift [23], sit/lie down [12], sit [17, 58, 33], grasp [52, 44]), or investigate a single type of object (e.g.
+chair [17, 58]).
 In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig.
 1.
-Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [tevet2023human] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt.
+Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [45] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt.
 To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt.
@@
 Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
-Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules.
-To our best knowledge, HOI-Diff is the first text-driven approach capable of generating realistic motions for various HOIs and different types of objects during the interactions.
-For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [bhatnagar22behave] with text descriptions, which bridges the gap of no data for text-driven 3D HOIs generation.
+Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
+To our best knowledge, HOI-Diff is one of the first text-driven approach s capable of generating realistic motions for various HOIs and different types of objects during the interactions.
+While our approach pioneers this capability, we note the existence of concurrent works such as  [9, 24, 49], which also explore similar directions.
+For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [3] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation.
+In addition, we evaluate our approach on the OMOMO dataset [25], which focuses on the manipulation of two hands.
 Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@
 Human Motion Generation with Diffusion Models.
-The denoising diffusion models have been widely used 2D image generations [rombach2022high, saharia2022photorealistic, ramesh2021zero] and achieved impressive results.
-Recent work [zhang2022motiondiffuse, tevet2023human, chen2023executing, karunratanakul2023gmd, rempeluo2023tracepace, ahn2023can, barquero2023belfusion, chen2023humanmac, dabral2023mofusion, shafir2023human, sun2023towards, tian2023transfusion, wei2023understanding, zhang2023generating, zhang2023remodiffuse, zhang2023tedi, xie2023omnicontrol] apply the diffusion model in the task of human motion generation.
+The denoising diffusion models have been widely used 2D image generations [39, 40, 36] and achieved impressive results.
+Recent work [56, 45, 5, 18, 38, 1, 2, 4, 6, 41, 43, 46, 51, 55, 57, 59, 53] apply the diffusion model in the task of human motion generation.
 While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with.
@@
 Scene- and Object-Aware Human Motion Generation.
-Recent works condition motion synthesis on scene geometry [huang2023diffusion, zhao2023synthesizing, wang2022towards, wang2022humanise].
+Recent works condition motion synthesis on scene geometry [15, 60, 48, 50].
 This facilitates the understanding of human-scene interactions.
 However, the motion fidelity is compromised due to the lack of paired full scene-motion data.
-Other approaches [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, Pi_2023_ICCV] instead focus on the interactions with the objects and can produce realistic motions.
+Other approaches [23, 17, 12, 42, 58, 33] instead focus on the interactions with the objects and can produce realistic motions.
 However, they focus on interacting with static objects with limited interactions.
-OMOMO [li2023object] proposes a conditional diffusion framework that can generate full-body motion from the object motion.
+OMOMO [25] proposes a conditional diffusion framework that can generate full-body motion from the object motion.
 The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion.
-IMoS [ghosh2023imos] synthesizes the full-body human along with the 3D object motions from textual inputs.
+IMoS [10] synthesizes the full-body human along with the 3D object motions from textual inputs.
 Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands.
-InterDiff [xu2023interdiff] predicts whole-body interactions with dynamic objects.
+InterDiff [54] predicts whole-body interactions with dynamic objects.
 However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information.
@@
 Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt.
+Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [9, 24, 49].
 Multi-Human Motion Generation.
-Recent works [shafir2023human, liang2023intergen, shafir2023human] incorporate human-to-human interactions into the motion diffusion process.
+Recent works [41, 26, 41] incorporate human-to-human interactions into the motion diffusion process.
 They usually also adopt cross-attention to exchange information from different human motions.
@@
 Affordance Estimation.
-The affordance estimation on 3D point cloud is studied in [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive].
-[kim2014semantic] formulates this task as the point cloud segmentation task from the extracted geometric features.
-[kim2015interactive] voxelizes the point cloud and creates affordance maps with interactive manipulation.
-[kokic2017affordance, iriondo2021affordance] focuses on the object grasping affordances and some specific applications.
-[mo2022o2o] learns the affordance for object-object interactions.
-[ngyen2023open] addressed the task of open-vocabulary affordance learning.
+The affordance estimation on 3D point cloud is studied in [30, 7, 22, 16, 29, 20, 21].
+[20] formulates this task as the point cloud segmentation task from the extracted geometric features.
+[21] voxelizes the point cloud and creates affordance maps with interactive manipulation.
+[22, 16] focuses on the object grasping affordances and some specific applications.
+[29] learns the affordance for object-object interactions.
+[30] addressed the task of open-vocabulary affordance learning.
 Overall affordance learning is a very challenging task even the shape of point cloud is given.
@@
 It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence.
-For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [Guo_2022_CVPR] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels.
+For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [11] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels.
 For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$.
 We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
-Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models.
-Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
 Diffusion Model for 3D HOI Generation.
-Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$***We use superscripts $h$ and $o$ to denote human and object squence, respectively.
+Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$ [^1] We use superscripts $h$ and $o$ to denote human and object sequence, respectively.
 Without a superscript, it means the 3D HOI sequence, containing both ${\bm{x}}^{h}$ and ${\bm{x}}^{o}$.
@@
 where $t$ is the denoising step.
-Following [tevet2023human], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
-We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [nichol2021improved]
+Following [45], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
+Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models.
+Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
+We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [31]
 ${\bm{\mu}}_{t}=\frac{\sqrt{\alpha_{t-1}}\beta_{t}}{1-\alpha_{t}}{\bm{x}}_{0}+\frac{\sqrt{1-\beta_{t}}(1-\alpha_{t-1})}{1-\alpha_{t}}{{\bm{x}}_{t}},$ (2)
 where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule.
-$\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [Ho2020-ew] is a variance scheduler of choice.
+$\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [14] is a variance scheduler of choice.
 Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
@@
 A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
-Simply adopting the diffusion model described in Eq.(1) would impose huge burden to the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description.
-In this paper, we propose HOI-Diff for 3D HOIs generation, which disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules.
+The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
+Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation.
+It has two branches designed for generating human and object motions individually.
+A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
+The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
+Figure 4: Illustration of APDM architecture for affordance estimation.
+Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean estimation.
+The object point cloud and textual prompt are taken as conditional input.
+Simply adopting the diffusion model described in Eq.(1) would impose a huge burden on the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description.
+In this paper, we propose HOI-Diff for 3D HOIs generation, disentangling motion generation for humans and objects and estimation of their contacting points.
+They are later integrated to form coherent and diverse HOIs, which reduces the complexity and burden for each of the three modules, leading to better generation performance as evidenced by our experiments.
 Fig.
@@
 As shown in Fig.
-3, it consists of two Transformer models [vaswani2017attention], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [tevet2023human].
-Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predicts clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
+4, it consists of two Transformer models [47], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [45].
+Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
 To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$.
 $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$.
-It then processes these inputs to generate refined updates based on the cross attention mechanism [vaswani2017attention].
+It then processes these inputs to generate refined updates based on the cross attention mechanism [47].
 The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
 The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
-Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [tevet2023human].
+Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [45].
 This fine-tuning is critical to ensure the smoothness of the generated human motions.
-We ablate this design choice in LABEL:ablation.
-Figure 4: Illustration of APDM architecture for affordance estimation.
-Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean affordance estimation.
-The object point cloud and textual prompt are taken as conditional input.
+We ablate this design choice in Sec.
+4.3.
+Object MDM is trained from scratch.
+We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion.
+More methodology details of HOI-DM can be found in the supplementary material.
 3.3 Affordance Estimation
 Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations.
-To improve the generation of intricate interactions, the first problem that needs to be solved is to identify where the contacting areas are between the human and object.
-InterDiff [xu2023interdiff] defines the contacting area based on the distance measurement between the surface of human and object.
+To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object.
+InterDiff [54] defines the contacting area based on the distance measurement between the surface of human and object.
 This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results.
@@
 As illustrated in Fig.
-4, the input to this model includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$.
+4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$.
 Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM.
 In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
-Affordance estimation in 3D point clouds itself is a notably challenging problem [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive], especially in the context of 3D HOI generation involving textual prompt.
-To address the limited availability of training data, we propose a simplified approach that diverges from predicting point-wise contact labels.
-Our method is grounded in two key hypotheses.
-Firstly, we posit that interactions with an object involve a maximum of two contact points.
-Secondly, we identify eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the primary interacting parts in most HOI scenarios.
-These hypotheses streamline the complexity of the problem, making it more tractable without significantly compromising accuracy.
-For instance, this approach effectively models common interactions such as grasping an object with both hands, sitting actions involving the pelvis, or lifting with a single hand.
-Our empirical observations suggest that these simplifications are valid in a wide range of HOI applications, providing a balance between model complexity and real-world applicability.
-Specifically, at each diffusion time step $n$ of APDMWe note that APDM and HOI-DM work independently.
-We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy affordance data ${\bm{y}}_{n}$ includes human contact labels $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, two object contact positions $\bm{y}_{n}^{o}\in\mathbb{R}^{2\times 3}$.
-In addition, our APDM also generates a binary object state label ${\bm{y}}_{n}^{s}\in\{0,1\}$ indicating whether it remains static or moves during the interaction.
-The generated binary label can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section.
+Affordance estimation in 3D point clouds itself is a notably challenging problem [30, 7, 22, 16, 29, 20, 21], especially in the context of 3D HOI generation involving textual prompt.
+In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios.
+It can effectively model common interactions such as grasping an object with both hands, sitting actions involving the pelvis and back, or lifting with a single hand.
+We use binary contact labels to determine which joints are in contact with the object.
+Subsequently, we predict eight corresponding contact points on the object surface, identified as the points closest to the selected body joints.
+Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
+Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and HOI-DM work independently.
+We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$.
+The model is designed to predict both contact probabilities and contact positions.
+Subsequently, dynamic selection of contacting body joints is performed by considering predicted probabilities over a specific threshold $\tau$ (set to be 0.6).
+The corresponding contact points on the object are then determined based on the selected joints.
 APDM works similar to the diffusion denoising process described in Eq.(1).
-Its output include clean affordance information ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
+Besides, we utilize a large language model (ChatGPT) to determine whether the object state ${\bm{y}}_{0}^{s}\in\{0,1\}$ should be set to static (${\bm{y}}_{0}^{s}=1$) based on the textual description, which can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section.
+All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
 3.4 Affordance-guided Interaction Correction
 With the estimated affordance, we can better align human and object motions to form coherent interactions.
-To this end, we propose to use the classifier guidance [dhariwal2021diffusion] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
-Algorithm 1 Affordance-guided Interaction Correction
-1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
-2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$
-3:$K$ = 1
-4:for all $t$ from $T$ to $1$ do
-5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$
-6:     Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$
-7:     if $t=1$ then
-8:         $K=100$
-9:     end if
-10:     for all $k$ from $K$ to $1$ do
-11:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$
-12:         ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$
-13:     end for
-14:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$
-15:     $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$
-16:end for
-17:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
+To this end, we propose to use the classifier guidance [8] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
+Algorithm 1 Affordance-guided Interaction Correction 1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
+2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$ 3:$K$ = 1 4:for all $t$ from $T$ to $1$ do 5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$ 6:     if $t=1$ then 7:         $K=100$ 8:     end if 9:     for all $k$ from $K$ to $1$ do # Separately perturb 10:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$ 11:     end for 12:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$ 13:end for 14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
 Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective.
 In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally.
-Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can purturb the generated human and object motion at each diffusion step $t$ as in [xie2023omnicontrol, karunratanakul2023guided, kulkarni2023nifty],
+Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [53, 19, 23],
 $\mathbf{{\bm{\mu}}}^{h}_{t}$ $=\mathbf{{\bm{\mu}}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}^{h}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}),$ (3)
@@
 We therefore minimize the distance between human contact joints and object contact points
-$G_{con}=\sum_{i\in\{1,2\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
-where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively.
-If there is just a single predicted contacting joint, $i=1$.
+$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
+where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th available contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively.
 $R(\cdot)$ converts the human joint’s local positions to global absolute locations, and $V(\cdot)$ obtains the object’s contact point sequence from the predicted mean of object pose ${\bm{\mu}}^{o}_{t}$.
@@
 However, we observe that when the human interacts with a static object, such as sitting on a chair, the object appears slightly moved.
-To address this, we utilize the predicted object states ${\bm{y}}_{0}^{s}$ to immobilize the object’s movement in the generated samples if the state is static, ensuring that proper contact is established between the human and the static object.
+To address this, we immobilize the object’s movement in the generated samples if the state is static (${\bm{y}}_{0}^{s}=1$), ensuring that proper contact is established between the human and the static object.
 The objective is defined as
-$\small{G_{sta}=\sum_{l=1}^{L}{\bm{y}}_{0}^{s}(l)\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
+$\small{G_{sta}={\bm{y}}_{0}^{s}\cdot\sum_{l=1}^{L}\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
 where ${\bm{\mu}}_{t}^{o}(l)$ denotes the object’s 6DoF pose in the $l$-th frame.
 $\bar{{\bm{\mu}}}_{t}^{o}=\frac{1}{L}\sum_{l}{\bm{\mu}}_{t}^{o}(l)$, which is the average of predicted means of the object’s pose.
-Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to space limit, we explain it in the supplementary material.
+Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in the supplementary material.
 Finally, we combine all these goal functions to as the final objective
@@
 4 Experiments
-Table 1: Quantitative results on the BEHAVE [bhatnagar22behave] dataset.
-We compare our method with three baselines adapted from existing models and the real data.
-The right arrow normal-$\rightarrow$ means closer to real data is better.
+4.1 Setup
+Dataset.
+Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [3].
+We then use ChatGPT to assist in generating three text descriptions for each HOI sequence, increasing the diversity of the data.
+Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects.
+It provides the human SMPL-H representation [27], the object mesh, as well as its 6DoF pose information in each HOI sequence.
+To ensure consistency in our approach, we follow the processing method used in HumanML3D [11] to extract representations for 22 body joints.
+All the models are trained to generate $L=196$ frames in our experiments.
+In the end, we have 1451 3D HOI sequences along with textual descriptions to train and evaluate our proposed approach.
+We follow the official train/test split on BEHAVE.
+We provide more details of the dataset, our annotation process, and annotated textual examples in the supplementary material.
+In addition, we evaluate our approach on OMOMO dataset [25].
+OMOMO focuses on full-body manipulation with hands.
+It consists of human-object interaction motion for 15 objects in daily life, with a total duration of approximately 10 hours.
+It provides text descriptions for each interaction motion.
+We utilize their object split strategy for both training and evaluation, ensuring the objects between the training and testing sets are different.
+Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset.
+More details of the OMOMO dataset are in the supplementary material.
+Evaluation metrics.
+We first assess different models for human motion generation using standard metrics as introduced by [11], namely Fréchet Inception Distance (FID), R-Precision, and Diversity.
+FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder.
+R-Precision gauges the relevance between generated motions and their corresponding text prompts.
+Diversity evaluates the range of variation in the generated motions.
+Additionally, we compute the Foot Skating Ratio to measure the proportion of frames exhibiting foot skid over a threshold (2.5 cm) during ground contact (foot height $<$ 5 cm).
+To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points.
+Ideally, we should develop similar metrics, e.g., FID, to evaluate the stochastic HOI generation.
+However, due to the limited data available in BEHAVE [3], training a motion encoder would produce biased evaluation results.
+To mitigate this issue, we resort to user studies to quantify the effectiveness of different models.
+Details will be introduced later.
+4.2 Comparisons with Existing Methods
+[…]
+[…]
 FID $\downarrow$
@@
 […]
-MDM* [tevet2023human] 3.169 0.206 7.606 0.448 0.184 1.20%
-PriorMDM* [shafir2023human] 4.352 0.215 8.186 0.416 0.270 21.42%
-Ours (Full) 1.632 0.222 7.596 0.363 0.190 77.38%
+MDM* [45] 3.169 0.206 7.606 0.448 0.184 0.87%
+PriorMDM* [41] 4.352 0.215 8.186 0.416 0.270 16.53%
+InterDiff [54] 7.192 0.170 7.686 0.506 0.218 0.00%
+Ours (Full) 1.501 0.235 7.391 0.347 0.182 82.60%
+Table 2: Quantitative results on OMOMO dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
+FID $\downarrow$
+R-precision $\uparrow$
+Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
+Real 1.539 0.193 4.833 - -
+MDM* 9.155 0.137 5.353 0.768 0.191
+PriorMDM 11.192 0.128 7.404 0.523 0.344
+InterDiff 10.567 0.110 6.493 0.906 0.239
+Ours 8.661 0.168 4.780 0.326 0.141
+Figure 5: Qualitative comparisons of our approach and baseline methods on BEHAVE dataset.
+The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4.
+This contrasts with the baselines, which fail to achieve a similar level of realism and contact plausibility in the interactions.
+As an additional visual aid, the mesh color gradually darkens over time to represent progression.
+(Best viewed in color.)
+Quantitative Results.
+Sec.
+4.2 reports the quantitative results on BEHAVE dataset [3].
+Compared with the baseline methods, our full method achieves the best performance.
+Specifically, it achieves state-of-the-art results in both FID, R-precision, and Diversity, underscoring its ability to generate high-quality human motions in the context of coherently interacting with objects.
+The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects.
+We also show the results of our User Study in Tab.
+4.2 to illustrate user preferences between our method and other baselines.
+In our user study, we sample 24 sequences from each method and 23 participants are asked to choose their most preferred generation results from these samples.
+This user study is designed to directly compare user preferences across different methods.
+The result indicates a strong preference for our method: it is favored over the baselines in a substantial majority of 82.60% of the cases.
+Tab.
+4.2 presents the quantitative results on the OMOMO dataset.
+Our method consistently outperforms other baselines by a considerable margin across all metrics.
+Notably, due to the distinctiveness of objects in the training and testing sets, the results in Tab.
+4.2 indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
 Table 3: Ablation studies of our model’s variants on the BEHAVE dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
-Supplementary Material
+FID $\downarrow$
+R-precision $\uparrow$
+Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
+User Study $\uparrow$
+w/o Interaction Correction
+Ours w/o CM 2.825 0.201 7.850 0.524 0.265 0.00%
+Ours w/o pretrain 2.414 0.226 7.915 0.402 0.158 3.90%
+[…]
+Ours 1.808 0.217 7.718 0.416 0.205 11.69%
+w/ Interaction Correction
+Ours w/o ${M}^{o}$ CM 3.206 0.205 7.909 0.365 0.310 0.00%
+Ours ${}^{Joint}$ 3.412 0.195 7.878 0.421 0.342 -
+Ours w/o $G_{con}$ 1.983 0.205 7.826 0.417 0.196 -
+Ours w/o $G_{sta}$ 1.636 0.213 7.785 0.367 0.181 -
+Ours w/o $G_{smo}$ 1.665 0.205 7.840 0.370 0.182 -
+Ours (Full) 1.501 0.235 7.391 0.347 0.182 84.41%
+Figure 6: Visual results of different variants of our model in ablation studies on the BEHAVE dataset [3].
+Qualitative Results.
+We showcase qualitative comparisons, rendered with SMPL [27] shapes, between our approach and the baseline methods in Fig.
+5.
+It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g., the toolbox in Fig.
+5 (b)).
+Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g., the chair in Fig.
+5 (e)).
+In stark contrast, our approach excels in creating visually appealing and realistic HOIs.
+Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g., raising the trash bin with the right hand in Fig.
+5 (a)).

### v3

@@
-https://neu-vi.github.io/HOI-Diff/
+HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models
 Abstract
@@
 To this end, we take a modular design and decompose the complex task into simpler sub-tasks.
-We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches.
+We first develop a dual-branch diffusion model (DBDM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches.
 We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt.
-The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter.
+The APDM is independent of the results by the DBDM and thus can correct potential errors by the latter.
 Moreover, it stochastically generates the contacting points to diversify the generated motions.
 Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects.
-To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions.
+To train and evaluate our approach, we annotate the BEHAVE dataset with text descriptions.
 Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
-Figure 1: HOI-Diff generates realistic motions for 3D human-object interactions given a text prompt and object geometry.
-Please see the sup.
-mat.
-for video results.
+Figure 1: HOI-Diff can generate realistic motions for 3D human-object interactions given a text prompt and object geometry.
+Please see the supplementary material for video results.
 Darker color indicates later frames in the sequence.
 Best viewed in color.
+The first two authors contributed equally.
 1 Introduction
@@
 First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects.
-Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair).
-Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored.
-Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [11, 34] or textual descriptions [3].
-On one hand, recent methods [23, 17, 12, 42, 58, 52, 44, 33] can synthesize realistic human motions for HOIs for static objects only.
-They usually synthesize the motion in the last mile of interaction, i.e.
-the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them.
+Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g, a person carries a chair, sits on a chair, pushes or pulls a chair).
+Text-driven 3D HOI synthesis with a diverse set of interactions is not yet fully addressed.
+Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either diverse HOIs [Guo_2022_CVPR, kit2016, li2023controllable] or detailed textual descriptions with interacting body parts and action [bhatnagar22behave, diller2023cghoi].
+It is important to note that CG-HOI [diller2023cghoi] has not made their code or annotations publicly available.
+In contrast, we will release both our code and annotations.
+Current methods cannot fully handle all the challenges.
+On one hand, recent methods [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, wu2022saga, taheri2021goal, Pi_2023_ICCV] can synthesize realistic human motions for HOIs for static objects only.
+They usually synthesize the motion in the last mile of interaction, i.e, the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them.
 On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity.
-For instance, they focus on grasping small objects [10], provide the object motion as conditioning [25], predict deterministic interactions between the human and the same object without the diversity [54, 37], consider only a small set of interactions (e.g.
-sit/lift [23], sit/lie down [12], sit [17, 58, 33], grasp [52, 44]), or investigate a single type of object (e.g.
-chair [17, 58]).
-In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig.
-1.
-Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [45] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt.
+For instance, they focus on grasping small objects [ghosh2023imos], provide the object motion as conditioning [li2023object], predict deterministic interactions between the human and the same object without the diversity [xu2023interdiff, razali2023action], consider only a small set of interactions (e.g., sit/lift [kulkarni2023nifty], sit/lie down [hassan_samp_2021], sit [jiang2022chairs, zhang2022couch, Pi_2023_ICCV], grasp [wu2022saga, taheri2021goal]), or investigate a single type of object (e.g., chair [jiang2022chairs, zhang2022couch]).
+In this paper, we introduce HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Figure 1.
+Our key insight is to decompose 3D HOIs synthesis into three modules to reduce the complexity of this challenging task.
+(a) coarse 3D HOIs generation that extends the human motion diffusion model [tevet2023human] to a dual-branch diffusion model (DBDM) to generate both human and object motions conditioning on the input text prompt.
 To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt.
-Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors.
+Our APDM does not rely on the results of the DBDM and thus can recover from its potential errors.
 Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
-Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
-To our best knowledge, HOI-Diff is one of the first text-driven approach s capable of generating realistic motions for various HOIs and different types of objects during the interactions.
-While our approach pioneers this capability, we note the existence of concurrent works such as  [9, 24, 49], which also explore similar directions.
-For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [3] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation.
-In addition, we evaluate our approach on the OMOMO dataset [25], which focuses on the manipulation of two hands.
+Compared with designing a monolithic model, HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
+For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [bhatnagar22behave] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation.
+In addition, we evaluate our approach on the OMOMO dataset [li2023object], which focuses on the manipulation of two hands.
 Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@
 Human Motion Generation with Diffusion Models.
-The denoising diffusion models have been widely used 2D image generations [39, 40, 36] and achieved impressive results.
-Recent work [56, 45, 5, 18, 38, 1, 2, 4, 6, 41, 43, 46, 51, 55, 57, 59, 53] apply the diffusion model in the task of human motion generation.
+The denoising diffusion models have been widely used 2D image generations [rombach2022high, saharia2022photorealistic, ramesh2021zero] and achieved impressive results.
+Recent work [zhang2022motiondiffuse, tevet2023human, chen2023executing, karunratanakul2023gmd, rempeluo2023tracepace, ahn2023can, barquero2023belfusion, chen2023humanmac, dabral2023mofusion, shafir2023human, sun2023towards, tian2023transfusion, wei2023understanding, zhang2023generating, zhang2023remodiffuse, zhang2023tedi, xie2023omnicontrol] apply the diffusion model in the task of human motion generation.
 While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with.
@@
 Scene- and Object-Aware Human Motion Generation.
-Recent works condition motion synthesis on scene geometry [15, 60, 48, 50].
+Recent works condition motion synthesis on scene geometry [huang2023diffusion, zhao2023synthesizing, wang2022towards, wang2022humanise].
 This facilitates the understanding of human-scene interactions.
 However, the motion fidelity is compromised due to the lack of paired full scene-motion data.
-Other approaches [23, 17, 12, 42, 58, 33] instead focus on the interactions with the objects and can produce realistic motions.
+Other approaches p[kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, Pi_2023_ICCV] instead focus on the interactions with the objects and can produce realistic motions.
 However, they focus on interacting with static objects with limited interactions.
-OMOMO [25] proposes a conditional diffusion framework that can generate full-body motion from the object motion.
+OMOMO [li2023object] can generate full-body motion from the object motion.
 The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion.
-IMoS [10] synthesizes the full-body human along with the 3D object motions from textual inputs.
-Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands.
-InterDiff [54] predicts whole-body interactions with dynamic objects.
-However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information.
-Note that the interaction type is deterministic, given the object and past motion information.
+IMoS [ghosh2023imos] synthesizes the full-body human along with the 3D object motions from textual inputs, but it only focuses on grasping small objects with hands.
+InterDiff [xu2023interdiff] predicts whole-body interactions with dynamic objects.
+Note that the interaction type is deterministic.
 Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt.
-Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [9, 24, 49].
-Multi-Human Motion Generation.
-Recent works [41, 26, 41] incorporate human-to-human interactions into the motion diffusion process.
-They usually also adopt cross-attention to exchange information from different human motions.
-Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
+Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [diller2023cghoi, wang2023physhoi, li2023controllable, song2024hoianimator, xu2024interdreamer].
+CG-HOI [diller2023cghoi] and HOIAnimator [song2024hoianimator] uses SMPL parameters as the motion representation, which may result in unsmooth motion due to the potential difficulty in optimization.
+Instead, we use common skeletal joints similar to most text-to-motion methods, harnessing the power of pre-trained human motion generation models.
+Chois li2023controllable relies on the initial state and object waypoints to generate HOIs, which reduces motion diversity for both the human and the object.
+InterFusion [dai2024interfusion] and F-HOI [yang2024f] generate static 3D HOIs from text description, lacking both human and object motions.
+Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models.
+Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (DBDM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) affordance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
 Affordance Estimation.
-The affordance estimation on 3D point cloud is studied in [30, 7, 22, 16, 29, 20, 21].
-[20] formulates this task as the point cloud segmentation task from the extracted geometric features.
-[21] voxelizes the point cloud and creates affordance maps with interactive manipulation.
-[22, 16] focuses on the object grasping affordances and some specific applications.
-[29] learns the affordance for object-object interactions.
-[30] addressed the task of open-vocabulary affordance learning.
-Overall affordance learning is a very challenging task even the shape of point cloud is given.
-Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
+The affordance estimation on 3D point cloud is studied in ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive.
+Overall affordance learning is a very challenging task.
+Instead of predicting the point-wise contact labels, we simplify it by directly regressing the contact points for human-object interactions, making it more tractable without significantly compromising accuracy.
 3 Method
-3.1 Overview
+The overview of our proposed approach are illustrated in Figure 2.
+We introduce a dual-branch Human-Object Interaction Diffusion Model (DBDM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec.
+3.2).
+To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec.
+3.3), whose output will be used as classifier guidance (Sec.
+3.4) to correct the interactions at each diffusion step of human/object motion generation.
+3.1 Background
 Motion Representations.
@@
 It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence.
-For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [11] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels.
-For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$.
+For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [Guo_2022_CVPR] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels.
+For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e, $D^{o}=6$.
 We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
@@
 where $t$ is the denoising step.
-Following [45], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
-Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models.
-Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
-We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [31]
+Following [tevet2023human], our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
+We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [nichol2021improved]
 ${\bm{\mu}}_{t}=\frac{\sqrt{\alpha_{t-1}}\beta_{t}}{1-\alpha_{t}}{\bm{x}}_{0}+\frac{\sqrt{1-\beta_{t}}(1-\alpha_{t-1})}{1-\alpha_{t}}{{\bm{x}}_{t}},$ (2)
 where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule.
-$\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [14] is a variance scheduler of choice.
+$\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [Ho2020-ew] is a variance scheduler of choice.
 Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
-Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation.
+Simply adopting the diffusion model described in Eq.(1) would impose a huge burden on the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description.
+In this paper, we propose HOI-Diff for 3D HOIs generation, disentangling motion generation for humans and objects and estimation of their contacting points.
+They are later integrated to form coherent and diverse HOIs, which reduces the complexity and burden for each of the three modules, leading to better generation performance as evidenced by our experiments.
+3.2 Coarse 3D HOIs Generation
+First, we introduce a dual-branch diffusion model (DBDM) to generate human and object motions that are roughly coherent.
+As shown in  Figure 4, it consists of two Transformer models [vaswani2017attention], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [tevet2023human].
+Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
+To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$.
+$CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$.
+It then processes these inputs to generate refined updates based on the cross attention mechanism [vaswani2017attention].
+The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
+The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
+Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [tevet2023human].
+This fine-tuning is critical to ensure the smoothness of the generated human motions.
+We ablate this design choice in Sec.
+LABEL:ablation.
+Object MDM is trained from scratch.
+We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion.
+More details of DBDM are in Appendix A.1.
+[…]
+Figure 3: Illustration of DBDM architecture for coarse 3D HOIs generation.
 It has two branches designed for generating human and object motions individually.
 A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
-The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
-Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation.
-It has two branches designed for generating human and object motions individually.
-A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
-The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
+The human motion model $M^{h}$ finetunes a pretrained MDM [tevet2023human].
+[…]
 Figure 4: Illustration of APDM architecture for affordance estimation.
@@
 The object point cloud and textual prompt are taken as conditional input.
-Simply adopting the diffusion model described in Eq.(1) would impose a huge burden on the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description.
-In this paper, we propose HOI-Diff for 3D HOIs generation, disentangling motion generation for humans and objects and estimation of their contacting points.
-They are later integrated to form coherent and diverse HOIs, which reduces the complexity and burden for each of the three modules, leading to better generation performance as evidenced by our experiments.
-Fig.
-2 shows the overview of our proposed approach.
-We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec.
-3.2).
-To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec.
-3.3), whose output will be used as classifier guidance (Sec.
-3.4) to correct the interactions at each diffusion step of human/object motion generation.
-3.2 Coarse 3D HOIs Generation
-First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent.
-As shown in Fig.
-4, it consists of two Transformer models [47], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [45].
-Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
-To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$.
-$CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$.
-It then processes these inputs to generate refined updates based on the cross attention mechanism [47].
-The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
-The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
-Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [45].
-This fine-tuning is critical to ensure the smoothness of the generated human motions.
-We ablate this design choice in Sec.
-4.3.
-Object MDM is trained from scratch.
-We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion.
-More methodology details of HOI-DM can be found in the supplementary material.
 3.3 Affordance Estimation
-Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations.
+Due to the complexity of the interactions between a human and object, DBDM alone usually fails to produce physically plausible results, leading to floating objects or penetrations.
 To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object.
-InterDiff [54] defines the contacting area based on the distance measurement between the surface of human and object.
+InterDiff [xu2023interdiff] defines the contacting area based on the distance measurement between the surface of human and object.
 This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results.
-In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table.
+In addition, the contact area is diverse even with the same object and interaction type, e.g, “sit” can happen on either side of a table.
 To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation.
-As illustrated in Fig.
-4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$.
-Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM.
+As illustrated in Figure 4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$.
+Our APDM doesn’t rely on the results of the DBDM and thus can recover from the potential errors in DBDM.
 In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
-Affordance estimation in 3D point clouds itself is a notably challenging problem [30, 7, 22, 16, 29, 20, 21], especially in the context of 3D HOI generation involving textual prompt.
+Affordance estimation in 3D point clouds itself is a notably challenging problem [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive], especially in the context of 3D HOI generation involving textual prompt.
 In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios.
@@
 Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
-Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and HOI-DM work independently.
+Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and DBDM work independently.
 We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$.
@@
 All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
+More implementation details are in Appendix A.2.
 3.4 Affordance-guided Interaction Correction
 With the estimated affordance, we can better align human and object motions to form coherent interactions.
-To this end, we propose to use the classifier guidance [8] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
-Algorithm 1 Affordance-guided Interaction Correction 1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
-2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$ 3:$K$ = 1 4:for all $t$ from $T$ to $1$ do 5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$ 6:     if $t=1$ then 7:         $K=100$ 8:     end if 9:     for all $k$ from $K$ to $1$ do # Separately perturb 10:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$ 11:     end for 12:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$ 13:end for 14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
+To this end, we propose to use the classifier guidance [dhariwal2021diffusion] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
 Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective.
 In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally.
-Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [53, 19, 23],
+Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [xie2023omnicontrol, karunratanakul2023guided],
 $\mathbf{{\bm{\mu}}}^{h}_{t}$ $=\mathbf{{\bm{\mu}}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}^{h}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}),$ (3)
@@
 This applies greater strength to perturb object motion, facilitating feasible corrections for contacting joints.
-During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step, as illustrated in Algorithm 1.
+During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step.
+The details are illustrated in Algorithm 1 of Appendix.
 How can we define the objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$?
@@
 We therefore minimize the distance between human contact joints and object contact points
-$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
+$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{\mu}}^{o}_{t},{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
 where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th available contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively.
@@
 $\bar{{\bm{\mu}}}_{t}^{o}=\frac{1}{L}\sum_{l}{\bm{\mu}}_{t}^{o}(l)$, which is the average of predicted means of the object’s pose.
-Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in the supplementary material.
+Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in Appendix A.3.
 Finally, we combine all these goal functions to as the final objective
-$G=G_{cont}+\alpha G_{sta}+\beta G_{smo},$ (7)
-where $\alpha$ and $\beta$ are weights for balance.
+$G=G_{con}+\alpha G_{sta}+\beta G_{smo},$ (7)
+where $\alpha=500$ and $\beta=100$ are weights for balance.
 4 Experiments
@@
 Dataset.
-Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [3].
-We then use ChatGPT to assist in generating three text descriptions for each HOI sequence, increasing the diversity of the data.
+Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [bhatnagar22behave].
+We then use GPT-3.5 [chatgpt] to rephrase and generate three text descriptions for each HOI sequence, increasing the diversity of the data.
 Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects.
-It provides the human SMPL-H representation [27], the object mesh, as well as its 6DoF pose information in each HOI sequence.
-To ensure consistency in our approach, we follow the processing method used in HumanML3D [11] to extract representations for 22 body joints.
+It provides the human SMPL-H representation [SMPL:2015], the object mesh, as well as its 6DoF pose information in each HOI sequence.
+To ensure consistency in our approach, we follow the processing method used in HumanML3D [Guo_2022_CVPR] to extract representations for 22 body joints.
 All the models are trained to generate $L=196$ frames in our experiments.
@@
 We follow the official train/test split on BEHAVE.
-We provide more details of the dataset, our annotation process, and annotated textual examples in the supplementary material.
-In addition, we evaluate our approach on OMOMO dataset [25].
+We provide more details of the dataset and annotation process in Appendix I.
+In addition, we evaluate our approach on OMOMO dataset [li2023object].
 OMOMO focuses on full-body manipulation with hands.
@@
 Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset.
-More details of the OMOMO dataset are in the supplementary material.
+More details are in Appendix J.
 Evaluation metrics.
-We first assess different models for human motion generation using standard metrics as introduced by [11], namely Fréchet Inception Distance (FID), R-Precision, and Diversity.
+We first assess different models for human motion generation using standard metrics as introduced by [Guo_2022_CVPR], namely Fréchet Inception Distance (FID), R-Precision, and Diversity.
 FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder.
@@
 To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points.
-Ideally, we should develop similar metrics, e.g., FID, to evaluate the stochastic HOI generation.
-However, due to the limited data available in BEHAVE [3], training a motion encoder would produce biased evaluation results.
+Ideally, we should develop similar metrics, e.g, FID, to evaluate the stochastic HOI generation.
+However, due to the limited data available in BEHAVE [bhatnagar22behave], training a motion encoder would produce biased evaluation results.
 To mitigate this issue, we resort to user studies to quantify the effectiveness of different models.
@@
 4.2 Comparisons with Existing Methods
-[…]
-[…]
-FID $\downarrow$
-R-precision $\uparrow$
-Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
-User Study $\uparrow$
-Real 0.096 0.324 7.244 - - -
-[…]
-MDM* [45] 3.169 0.206 7.606 0.448 0.184 0.87%
-PriorMDM* [41] 4.352 0.215 8.186 0.416 0.270 16.53%
-InterDiff [54] 7.192 0.170 7.686 0.506 0.218 0.00%
-Ours (Full) 1.501 0.235 7.391 0.347 0.182 82.60%
-Table 2: Quantitative results on OMOMO dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
-FID $\downarrow$
-R-precision $\uparrow$
-Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
-Real 1.539 0.193 4.833 - -
-MDM* 9.155 0.137 5.353 0.768 0.191
-PriorMDM 11.192 0.128 7.404 0.523 0.344
-InterDiff 10.567 0.110 6.493 0.906 0.239
-Ours 8.661 0.168 4.780 0.326 0.141
-Figure 5: Qualitative comparisons of our approach and baseline methods on BEHAVE dataset.
+Baselines.
+Our work introduces a novel 3D HOIs generation task not addressed by existing text-to-motion methods, which focus exclusively on human motion generation without accounting for human-object interactions.
+To compare with existing works, we mainly focus on evaluating human motion generation.
+We then design different variants of our models for comparing 3D HOIs generation.
+Specifically, we adopt the prominent text-to-motion methods MDM [tevet2023human] and PriorMDM* [shafir2023human] with the following settings.
+(a) MDM: In this setup, we finetune the original MDM model [tevet2023human] on the BEHAVE dataset [bhatnagar22behave] without object motion.
+(b) MDM*: This variant involves adapting the input and output layers’ dimensions of the MDM model [tevet2023human] to accommodate the input of 3D HOI sequences.
+This adjustment allows for the simultaneous learning of both human and object motions within a singular, integrated model.
+(c) PriorMDM* [shafir2023human]: We adapt the ComMDM architecture proposed in [shafir2023human], originally designed for two-person motion generation, to suit our needs for HOIs synthesis by modifying one of its two branches for object motion generation.
+(d) InterDiff [xu2023interdiff]: While InterDiff is not designed for text-driven synthesis of 3D HOI, we added text conditioning to InterDiff as the baseline.
+More details are in Appendix C.
+BEHAVE OMOMO
+Method FID
+R-precision
+Diversity
+Contact
+Pene
+Foot Skate
+FID
+R-precision
+Diversity
+Contact
+Pene
+Foot Skate
+Ratio $\downarrow$
+Real 0.04 0.86 12.48 - - - 0.57 0.63 9.98 - - -
+MDM 6.77 0.34 10.81 - - - 12.28 0.23 5.56 - - -
+MDM* 4.25 0.38 11.23 0.448 0.52 0.190 10.37 0.21 6.04 0.768 0.41 0.191
+PriorMDM* 4.54 0.30 10.03 0.416 0.57 0.270 9.87 0.25 6.34 0.523 0.38 0.344
+InterDiff 8.58 0.26 10.75 0.506 0.42 0.218 14.27 0.17 5.69 0.906 0.32 0.239
+Ours 1.62 0.46 12.02 0.347 0.51 0.182 8.76 0.31 8.13 0.326 0.39 0.141
+Table 1: Quantitative results on the BEHAVE and OMOMO dataset.
+We compare our method with baselines adapted from existing models.
+MDM: fine-tune the original MDM [tevet2023human] on the BEHAVE dataset without object motion.
+MDM*: adapting the input and output layers’ dimensions of the MDM to accommodate both human and object motions.
+PriorMDM*: We adapt the ComMDM architecture proposed in shafir2023human.
+InterDiff: We add a CLIP encoder in xu2023interdiff to support our task.
+The right arrow $\rightarrow$ means closer to real data is better.
+Figure 5: Qualitative comparisons of our approach and baselines on BEHAVE dataset.
 The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4.
@@
 Quantitative Results.
-Sec.
-4.2 reports the quantitative results on BEHAVE dataset [3].
+Table 4.2-left reports the quantitative results on BEHAVE dataset [bhatnagar22behave].
 Compared with the baseline methods, our full method achieves the best performance.
@@
 The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects.
-We also show the results of our User Study in Tab.
-4.2 to illustrate user preferences between our method and other baselines.
-In our user study, we sample 24 sequences from each method and 23 participants are asked to choose their most preferred generation results from these samples.
-This user study is designed to directly compare user preferences across different methods.
-The result indicates a strong preference for our method: it is favored over the baselines in a substantial majority of 82.60% of the cases.
-Tab.
-4.2 presents the quantitative results on the OMOMO dataset.
+Table 4.2-right presents the quantitative results on the OMOMO dataset.
+We used the train/test split of the OMOMO dataset to evaluate the model’s inference capacity on unseen objects, including the small table, white chair, suitcase, and tripod.
 Our method consistently outperforms other baselines by a considerable margin across all metrics.
-Notably, due to the distinctiveness of objects in the training and testing sets, the results in Tab.
-4.2 indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
-Table 3: Ablation studies of our model’s variants on the BEHAVE dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
-FID $\downarrow$
-R-precision $\uparrow$
-Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
-User Study $\uparrow$
-w/o Interaction Correction
-Ours w/o CM 2.825 0.201 7.850 0.524 0.265 0.00%
-Ours w/o pretrain 2.414 0.226 7.915 0.402 0.158 3.90%
+Notably, due to the distinctiveness of objects in the training and testing sets, the results indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
+We also provide user study results, please refer to Appendix G for details.
+Qualitative Results.
+We showcase qualitative comparisons, rendered with SMPL [SMPL:2015] shapes, between our approach and the baseline methods in Figure 5.
+It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g, the toolbox in Figure 5 (b)).
+Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g, the chair in Figure 5 (e)).
+In stark contrast, our approach excels in creating visually appealing and realistic HOIs.
+Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g, raising the trash bin with the right hand in Figure 5 (a)).
+For the same object, our method can generate diverse HOIs using different body parts and contact points, as shown in Figure 16 in Appendix.
+Supplementary Material
+Appendix A Additional Details of Methodology
+In Sec.
+3 of our main paper, we presented the foundational design of each key component in our HOI-Diff pipeline.
+Here, we delve into an elaborate explanation of model architecture, learning objectives and additional details associated with each crucial component.
+A.1 Dual-branch diffusion model (DBDM)
+The Communication Module (CM) in DBDM is based on the cross attention mechanism. Formally,
+$\tilde{\bm{f}}^{h}=\operatorname{MLP}(\operatorname{Attn}({\bm{f}}^{h}{\mathbf{W}}_{Q},{\bm{f}}^{o}{\mathbf{W}}_{K},{\bm{f}}^{o}{\mathbf{W}}_{V})),$ (8)
+$\tilde{\bm{f}}^{o}=\operatorname{MLP}(\operatorname{Attn}({\bm{f}}^{o}{\mathbf{W}}_{Q},{\bm{f}}^{h}{\mathbf{W}}_{K},{\bm{f}}^{h}{\mathbf{W}}_{V})),$ (9)
+where $\operatorname{MLP}(\cdot)$ denotes fully-connected layers, $\operatorname{Attn}(\cdot)$ is the attention block [vaswani2017attention], and ${\mathbf{W}}_{Q},{\mathbf{W}}_{K},{\mathbf{W}}_{V}$ are learned projection matrices for query, key, and value, respectively.
+The training objective of this full model is based on reconstruction loss
+$\begin{split}\mathcal{L}_{hoi}=\mathbb{E}_{t\sim[1,T]}\|M_{\theta}(\bm{x}_{t},t,\bm{c})-\bm{x}_{0}\|^{2}_{2},\end{split}$ (10)
+where $\bm{x}_{0}$ is the ground truth of the HOI sequence.
+A.2 Affordance prediction diffusion model (APDM).
+Model architecture.
+The affordance prediction diffusion model comprises eight Transformer layers for the encoder with a PointNet++ [qi2017pointnet++] to encode the object’s point clouds.
+The training objective of this diffusion model is also based on reconstruction loss
+$\begin{split}\mathcal{L}_{aff}=\mathbb{E}_{t\sim[1,T]}\|{A}_{\theta}(\bm{y}_{t},t,\bm{p},\bm{d})-\bm{y}_{0}\|^{2}_{2},\end{split}$ (11)
+where $\bm{y}_{0}$ is the ground-truth affordance data.
+$\bm{p}$ and $\bm{d}$ denote object point cloud and text description (prompt), respectively.
+$A_{\theta}$ represents the affordance prediction diffusion model.
+Inferring object state with GPT-3.5-turbo in APDM.
+To infer the state of an object, we directly leverage the strong prior knowledge of large language models to derive the result.
+Specifically, we utilize the GPT-3.5-turbo [chatgpt] API by inputting specific instructions, allowing it to infer the result directly based on the input HOI text description.
+The prompt template for instruction is shown in Figure 6.
+A.3 Affordance-guided interaction correction.
+During the inference stage, it’s found that the predicted object contact positions may occasionally be inaccurately positioned, residing either inside or outside the object.
+To rectify this, we implement post-processing steps that replace these predicted contact points, denoted as $\bm{y}_{0}^{o}$, with their nearest neighbors from the object’s point clouds.
+This adjustment aims to enhance the accuracy of the updated contact points, aligning them more closely with their actual positions on the object’s surface.
+However, employing these updated contact points directly for contact constraints, particularly in the absence of detailed human shape information, introduces a new challenge.
+It can potentially lead to penetration issues within the contact area while reconstructing the human mesh in the final stage.
+To mitigate contact penetration, we adopt a method that recalculates points at a specified distance outward, perpendicular to the normal, originating from the object’s contact points.
+This process can formulated as: $\tilde{\bm{y}}{0}^{o}=\hat{\bm{y}}_{0}^{o}+v_{n}^{i}*d$, where $i\in\{1,2\}$ indicates the $i^{th}$ object contact points, $v_{n}^{i}$ denotes the normal vector at that point and $d=0.05$ is a contact distance threshold.
+As for smoothness term, we formulate it as
+$\begin{split}\small{G_{smo}=\sum_{l=1}^{L-1}\left\lVert{\bm{x}}_{0}^{o}(l+1)-{\bm{x}}_{0}^{o}(l)\right\rVert^{2},}\end{split}$ (12)
+where ${\bm{x}}_{0}^{o}(l)$ is the predicted 6DoF pose of the object in the $l$-th frame.
+Figure 6: Prompt template for inferring object state.
+Algorithm 1 Affordance-guided Interaction Correction
+1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
+2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$
+3:$K$ = 1
+4:for all $t$ from $T$ to $1$ do
+5:       ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$
+6:       if $t=1$ then
+7:             $K=100$
+8:       end if
+9:       for all $k$ from $K$ to $1$ do # Separately perturb
+10:             ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$
+11:       end for
+12:       $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$
+13:end for
+14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
+Appendix B Implementation Details
+Both our DBDM and APDM are built on the Transformer [vaswani2017attention] architecture.
+Similar to MDM [tevet2023human], we employ the CLIP model to encode text prompts, adhering to a classifier-free generation process.
+Our models are trained using PyTorch [paszke2019pytorch] on 1 NVIDIA A5000 GPU.
+We set control strength of guidance as $\tau_{1}$ = 1, $\tau_{2}$ = 100, and ${\Sigma}_{t}=min(\Sigma_{t},0.01)$.
+Both the DBDM and APDM are trained on the same data for 20k steps.
+Both the DBDM and APDM architectures of HOI-Diff are based on Transformers with 4 attention heads, a latent dimension of 512, a dropout of 0.1, a feed-forward size of 1024, and the GeLU activation [hendrycks2016gaussian].
+The number of learned parameters for each model is stated in Table 9.
+Our training setting involves 20k iterations for the DBDM and 10k iterations for the APDM model.
+These iterations utilize a batch size of 32 and employ the AdamW optimizer [loshchilov2017decoupled] with a learning rate set at 10-4.
+We use $T$=1000 and $N$=500 diffusion steps in DBDM and APDM, respectively.
+Appendix C Additional Details of Baselines
+  • MDMfinetuned: We finetune MDM [tevet2023human] on BEHAVE dataset without considering the object motion.
+  • MDM*: We extend the original feature dimensions of the input and output processing in MDM [tevet2023human] from $D^{h}$ to $D^{h}+D^{o}$, enabling support for HOIs sequences.
+The model is trained from scratch on BEHAVE dataset [bhatnagar22behave].
+  • PriorMDM*: The proposed approach for dual-person motion generation employs paired fixed MDMs [tevet2023human] per individual to ensure uniformity within generated human motion distributions.
+This design leverages a singular ComMDM to coordinate between the two branches of fixed MDM instances, streamlining training and maintaining consistency across generated motions.
+Given that both branches are based on MDM that pretrained on human motion datasets, direct utilization of them for human-object interactions in our task is infeasible.
+We maintain one branch dedicated to humans, leveraging pre-trained weights, while adapting the input and output processing of another branch specifically for generating object motion.
+Following this, we fine-tune the human MDM branch while initiating the learning of object motion from scratch within the object branch.
+Eventually, we integrate ComMDM to facilitate communication and coordination between these distinct branches handling human and object interactions.
+  • InterDiff: InterDiff [xu2023interdiff] is originally designed for a prediction task rather than text-driven HOIs generation.
+To tailor it to our task, we replace its Transformer encoder with a CLIP encoder and modify its feature dimensions of the input and output layers.
+To ensure fair comparisons, all the above baselines as well as our own models are all trained on BEHAVE and OMOMO datasets for 20k steps.
+Figure 7: Effect of different total numbers of perturbations in the whole denoising process.
+(a) Perturb one time in each denoising step (in total $T=1000$).
+(b) Perturb one time in first $T-1$ denoising steps, and repeatedly perturb 10 times in the final denoising step.
+(c) Perturb one time in first $T-1$ denoising steps, and repeatedly perturb 100 times in the final denoising step.
+Appendix D Additional Details of Evaluation Metrics
+For detailed information regarding metrics employed in human motion generation, including FID, R-Precision, and Diversity, we refer readers to [tevet2023human, Guo_2022_CVPR] for comprehensive understanding.
+Contact Distance.
+Expanding on the concept of Contact Distance, we utilize the chamfer distance metric to quantify the closeness between human body joints and the object surface.
+This computation leverages ground-truth affordance data that includes human contact labels and object contact points,
+$ContactDistance=\frac{1}{L}\sum_{l}^{L}{CD(\hat{\bm{x}}_{l}^{h},\hat{\bm{p}}_{l})},$ (13)
+where $\hat{\bm{x}}_{l}^{h}$ represents two human contact joints at the $l$-th frame, indexed according to ground-truth contact labels.
+Additionally, $\hat{\bm{p}}_{l}$ denotes two object contact points derived from the object motion $\bm{x}^{o}_{l}$ at frame $l$, also indexed based on ground-truth information.
+$CD$ denotes the chamfer distance.
+Penetration Score.
+We followed the  li2023controllable to compute the penetration score (Pene), each vertex of the body ($V_{i}$) is queried against the precomputed Signed Distance Field (SDF) of the object.
+This process yields a corresponding distance value for each vertex.
+The penetration score is then formalized as:
+$Pene=\frac{1}{n}\sum_{i=1}^{n}{|min(d_{i},0)|},$ (14)
+measured in centimeters (cm).
 […]
-Ours 1.808 0.217 7.718 0.416 0.205 11.69%
-w/ Interaction Correction
-Ours w/o ${M}^{o}$ CM 3.206 0.205 7.909 0.365 0.310 0.00%
-Ours ${}^{Joint}$ 3.412 0.195 7.878 0.421 0.342 -
-Ours w/o $G_{con}$ 1.983 0.205 7.826 0.417 0.196 -
-Ours w/o $G_{sta}$ 1.636 0.213 7.785 0.367 0.181 -
-Ours w/o $G_{smo}$ 1.665 0.205 7.840 0.370 0.182 -
-Ours (Full) 1.501 0.235 7.391 0.347 0.182 84.41%
-Figure 6: Visual results of different variants of our model in ablation studies on the BEHAVE dataset [3].
-Qualitative Results.
-We showcase qualitative comparisons, rendered with SMPL [27] shapes, between our approach and the baseline methods in Fig.
-5.
-It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g., the toolbox in Fig.
-5 (b)).
-Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g., the chair in Fig.
-5 (e)).
-In stark contrast, our approach excels in creating visually appealing and realistic HOIs.
-Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g., raising the trash bin with the right hand in Fig.
-5 (a)).
+capbtabboxtable[][]
+[…]
+Model DBDM APDM
+Parameters (·106) 8.82 38.92
+Figure 8: Model Parameters. The number of learned parameters of our two core architectures.
+[…]
+Method MDM* PriorMDM* Ours (Full)
+Time (s) 32.3 38.6 118.0
+Component APDM DBDM Interaction Correction
+Time (s) 24.2 46.4 47.4
+Figure 9: Inference Time (on NVIDIA A5000 GPU). We report the inference time for baselines, our full method, and its key components.
+Appendix E Inference Time
+In Table 9, we provide the inference times for both baselines and our full method, including its key components.
+All measurements were conducted using an NVIDIA A5000 GPU.
+Training an additional model for affordance information and using classifier guidance for interaction correction do contribute to increased inference costs.
+However, despite the longer inference time, our complete method notably enhances the accuracy of 3D HOIs generation.
+Params (M) FID $\downarrow$ R-precision (Top-3) $\uparrow$
+MDM 49.85 6.98 0.36
+Ours (Full) 47.74 1.62 0.46
+Table 4: With comparable model size, the performance results of MDM and Ours (Full).
+Figure 10: Effect of different control strengths for classifier guidance.
+(a) We use equal strengths of $\tau_{1}=1,\tau_{2}=1$ to perturb the predicted mean of human motion and object motion, respectively.
+(b) We use different strengths of $\tau_{1}=1,\tau_{2}=100$ for the perturbation.
+We can see that different strengths work better.
+Appendix F Additional Ablation Studies
+Different perturbing times in classifier guidance.
+As discussed in Sec.
+3.4, in the later stage of classifier guidance, diffusion models tend to strongly attenuate the introduced signals.
+Therefore, we iteratively perturb the predicted mean of motion for $K$ times at the final denoising step.
+In Figure 7, we present the ablation results, illustrating the impact of different numbers of perturbations.
+Notably, we observe that employing 100 perturbations leads to re-convergence and yields the desired results.
+Different guidance strength.
+As detailed in Sec.
+3.4, we employ distinct control strengths for classifier guidance, considering the varying feature densities in predicted human and object motion.
+Rather than employing equal control strengths, we opt to assign a higher control strength to object motion, allowing it to closely align with human contact joints, as illustrated in Figure 10.
+Different model with comparable model size.
+Although our method involves a slightly larger number of model parameters, our model is specifically designed for HOI generation.
+As seen in the  Table 4, if we attempt to scale MDM* to the same model size, its performance remains subpar.
+Appendix G User Study
+For each method, we select 15 prompts from the BEHAVE dataset and 10 prompts from the OMOMO dataset, covering various interaction types and object items.
+We sample twice with each prompt to gather a total of 50 results.
+40 participants are asked to choose their most preferred generation results from these samples.
+This user study requires pairwise comparisons of our method with other baseline on generated interaction quality, as shown in Figure 17.
+The results in Figure 12 indicate strong preference for our method: it is favored over the baselines in 89.6% (Ours vs.
+MDM*), 73.8% (Ours vs.
+PriorMDM*) and 95.3% (Ours vs.
+Interdiff).
+[…]
+Figure 11: Perceptual User Study. Most participants prefer our method over the baselines.
+[…]
+Figure 12: Visual results of different variants of our model in ablation studies.
+Appendix H Additional Qualitative Results
+In this section, we present additional qualitative results showcasing the model’s performance evaluated on the OMOMO dataset, and the effectiveness of APDM.
+Figure 13: Additional qualitative evaluation on OMOMO dataset.
+Given object geometry and text description, our method can generate high-quality human-object interactions even for the unseen objects (tripod, smalltable, suitcase).
+Figure 14: Additional qualitative evaluation on unseen objects.
+Figure 15: Visual results of estimated contact points.
+Our APDM, trained on the BEHAVE dataset, can accurately estimating contact positions for objects based on textual descriptions.
+Furthermore, it showcases the capability to generalize to unseen objects in the OMOMO dataset, as demonstrated in the last row.
+Qualitative results on OMOMO dataset.
+We present additional qualitative results on the OMOMO dataset, rendered with SMPL [SMPL:2015] shapes, as shown in Figure 13.
+It is evident that our method can generalizes effectively to unseen objects and produce realistic 3D human-object interactions.
+Qualitative results of APDM.
+To verify the accuracy of estimated contact points on object surface, we provide additional visual results in Figure 15.
+It can be seen that our method can predict realistic and practical contact points based on text descriptions.
+With APDM, we even can generate different interactions with the same object based on the input description, as shown in the Figure 16.
+Generalization capability.
+To verify the model’s generalization capability, except of unseen object test on OMOMO dataset, we also downloaded several objects from Sketchfab [^3] https://sketchfab.com/, adjusted them to a reasonable scale, and used them as inputs.
+As shown in Figure 14, our model successfully establishes reasonable HOI contact with these previously unseen objects.
+Figure 16: Leveraging the power of the APDM module, our method can generate diverse HOIs for the same object using different contacting body parts and contact points.
+Figure 17: An example question for our text-to-hoi user study.
+Appendix I Annotation for BEHAVE Dataset
+Text Annotating Process.
+Initially, we manually annotate the interaction types and the specific human body parts involved, delineating actions like “lift” associated with the “left hand” or “hold” involving “two hands”.
+Subsequently, to generate complete sentences, we leverage the capabilities of GPT-3.5 to assist in formulating the entirety of the description.
+Table 5: Examples of our annotated textual descriptions for the BEHAVE dataset rephrased by GPT-3.5 [chatgpt].
+Object Textual Descriptions
+backpack A person is carrying the backpack in front.
+The person is raising a backpack with his right hand.
+The person at the front presently has control over the backpack.
+chairwood A person is using the chairwood for sitting.
+(wooden chair) The person is propelling the chairwood on the ground.
+Someone is hoisting a chairwood by his left hand.
+tablesquare A person is lifting the tablesquare, utilizing his left hand.
+(square table) Someone is clutching onto a tablesquare from the front.
+An individual is moving the tablesquare back and forth.
+boxlong A person is gripping the boxlong from the front.
+(long box) A person is raising the boxlong using his left hand.
+Someone hoists the boxlong with his left hand.
+toolbox Someone is grasping the toolbox upfront.
+The person has a firm hold on the toolbox with his right hand.
+A person is gripping the toolbox with his left hand.
+yogaball A person is shifting a yogaball back and forth on the floor using his hands.
+The person is occupying a yogaball.
+A person is employing an yogaball to engage in an upper body game.
+Examples of Annotated Textual Descriptions. In Table 5, we showcase a selection of our annotated textual descriptions for the BEHAVE dataset [bhatnagar22behave].
+Figure 18: Analysis of word frequency We count the occurrences of each interaction verb from all text descriptions to illustrate their respective frequencies.
+Analysis of Annotated Textual Descriptions.
+All text descriptions encompass 36 distinct interaction verbs associated with 20 different objects.
+Figure 18 illustrates the frequency of each verb, indicating their respective occurrences.
+Affordance Data.
+Our affordance data includes 8-dimensional human contact labels and object contact points.
+We employ chamfer distance to measure the distance between all human body joints and object surface points.
+Following a predefined distance threshold $\gamma=0.12$, we identify the 8 contact points on the object surface corresponding to the 8 primary human body joints.
+Subsequently, we derive the human contact labels by encoding the indexes of contact joints into an 8-dimensional vector represented by binary values.
+Appendix J Additional Details of OMOMO Dataset
+The OMOMO dataset comprises data captured for a total of 15 objects.
+Adhering to their official split strategy depicted in [li2023object](Figure 5), we allocate 10 objects for training and 5 objects for testing.
+This split allows us to further evaluate the model’s generalization ability to new objects.
+Notably, the OMOMO dataset itself provides text annotation, and we use GPT-3.5 to add subjects to it and embellish it appropriately.
+For affordance data, we preprocess it the same way we handle BEHAVE.
+Appendix K Common Questions
+Why use Skeletal Pose Representation rather than SMPL parameters?
+Most state-of-the-art text-to-motion methods adopt the skeletal pose representation proposed by Guo_2022_CVPR, demonstrating excellent performance and stability.
+While some works [azadi2023make] argue that SMPL parameters [SMPL:2015] contains shape and global information, it does not generate as smooth motions as skeletal-based approaches.
+Consequently, we adopt the skeletal pose representation and aim to leverage strong pose priors from the pretrained text-to-motion model [tevet2023human] to ensure the authenticity of generated human motion.
+Can we handle multi-phase interactions between humans and objects?
+Due to the lack of fine-grained textural descriptions in the current 3D HOI dataset, we primarily consider only one interaction phase.
+However, we have found that an LLM can still reason well for multiple phases given a template such as: You will be given a sentence that describes an interaction between a person and an object across multiple phases.
+Your task is to divide the interaction into phases based on the state of the object and determine the state for each phase.
+If the object is being moved by the person during a phase, output the number 0.
+If the object remains stationary during a phase, output the number 1.
+For example, given the text description: The box is on the ground.
+A person is picking up the box and holding it forward, then putting the box towards the table.
+The box is on the table" The result from GPT-3.5-turbo: "Phase 1: The box is on the ground - State: 1 (stationary); Phase 2: The person is picking up the box and holding it forward - State: 0 (moved); Phase 3: The person is putting the box towards the table - State: 0 (moved); Phase 4: The box is on the table - State: 1 (stationary).
+We will address the generation of multiple phases of 3D HOI in future work.
+Can we generate hand motion with articulated fingers?
+The BEHAVE and OMOMO datasets do not capture and provide raw hand parameters, despite utilizing SMPLH and SMPLX models to fit human body meshes for rendering.
+Consequently, in this paper, we focus solely on whole-body human motion, excluding articulated hand and finger movements.
+Why do we use large language models (LLMs) to predict object state based on the input description?
+We aim to leverage LLMs for inferring object states, and our results demonstrate that they perform efficiently and effectively.
+As shown in the Table 6, we evaluated the performance of object state prediction with GPT-3.5-turbo [chatgpt] and obtained an average precision of 95.6% on the validation set, with an average response time of 0.518 seconds.
+The results suggest that GPT-3.5-turbo is sufficiently accurate without adding significant overhead.
+We also evaluated the prediction performance using other LLMs, including Gemini-1.5-Pro-Exp-0801 [reid2024gemini] (99.4%, 0.259s), Gemma-2-27B [team2024gemma] (98.6%, 0.522s), and LLaMA-2-13B [touvron2023llama] (94.4%, 0.521s), the latter two being publicly available.
+Acc (%) $\uparrow$ Time (s) $\downarrow$
+GPT-3.5 95.6 0.518
+Gemini-1.5-Pro-Exp-0801 99.4 0.259
+Gemma-2-27B 98.6 0.522
+LLaMA-2-13B 99.4 0.259
+APDM + MLP 79.5 2.420
+Table 6: LLMs’ inference accuracy (Acc) and average inference time (Time) on object state prediction.
+To further validate the effectiveness of the LLM module, we modified the APDM module by adding an MLP head to predict the object status.
+The newly added MLP takes in the features consisting of object geometry information and CLIP embeddings.
+We used an MSE loss.
+We got average precision 79.5% and average time 2.42s for this design on the validation set, which is significantly worse than the results of GPT-3.5-turbo (95.6%, 0.518s), Gemma-2-27b (98.6%, 0.522s), Gemini-1.5-Pro-Exp-0801 (99.4%, 0.259s) and LLaMA-2-13B (4.4%, 0.521s).
+In future work, we believe the LLM can play a more important role in 3D HOI, e.g.
+providing high-level instruction for more complex human-object interactions, and our initial use of the LLM offers insights into its potential applications and how it can be effectively utilized.
+Appendix L Supplementary Video
+Beyond the qualitative results presented in the main paper, our supplementary materials offer comprehensive demos that provide an in-depth visualization of our task, further showcasing the effectiveness of our approach.
+In these demonstrations, we highlight the better performance of our method, HOI-Diff, in producing diverse and realistic 3D HOIs while maintaining adherence to physical validity.
+Notably, the visualizations show that HOI-Diff consistently generates smooth, vivid interactions, accurately capturing human-object contacts.
+Additionally, we present the visual ablation results and emphasize the significance and effectiveness of our affordance-guided interaction correction, underscoring its substantial impact on improving the overall performance and quality of the generated 3D HOIs.
+Appendix M Limitations
+The existing datasets for 3D HOIs are limited in terms of action and motion diversity, posing a challenge for synthesizing long-term interactions in our task.
+Furthermore, the effectiveness of our model’s interaction correction component is contingent on the precision of affordance estimation.
+Despite simplifying this task, achieving accurate affordance estimation remains a significant challenge, impacting the overall performance of our model.
+A promising direction for future research involves integrating a sophisticated affordance model pre-trained on an extensive 3D object dataset, along with text prompts.
+Such an advancement could significantly enhance the realism and accuracy of human-object contact in our model, leading to more natural and precise HOIs synthesis.
+Appendix N Social Impacts
+On the positive side, it may offers the research community valuable insights into understanding human behaviors.
+On the negative side, it remains uncertain whether individuals can be identified solely based on their poses and movements.
+However, compared to traditional input images of people, this method poses a lower risk of invading personal privacy.